The point price elasticity of demand for red herring is -4. The demand curve for red herring is Q = 120-P.
What is the price of red herring?
1
Expert's answer
2016-04-04T09:21:04-0400
If the point price elasticity of demand for red herring is -4 and demand curve for red herring is Q = 120-P, then the formula for point price elasticity is Ed = Qd'*P/Q, so:
-4 = (-1)*P/Q,
P/Q = 4,
Q = P/4
If Qd = 120 - P, then:
120 - P = P/4,
1.25P = 120,
P = $96.+
So, the price of red herring is $96.